OpenStudy (anonymous):

Given that y = 1 when x = 0, solve the differential equation dy dx = y3 + 1 y2 , obtaining an expression for y in terms of x.

OpenStudy (anonymous):

sorry it is dy/dx = (y^3 +1)/y^2

OpenStudy (anonymous):

\[dy/dx = \frac{(y^{3} + 1)} {y^{2}}\] is that it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lol i feel silly but it's a no solution XD look at Dy/Dx X being zero, you can't divide by zero :P

OpenStudy (anonymous):

[y^2 / (y^3+1)] dy = dx integrating both sides y^3 + 1 = u du = 3y^2dy du/3 = y^2dy du / (3u) = dx x +C = (1/3) ln|u| x +C = (1/3) ln|y^3 + 1| 3e^(x+c) -1 = y^3 y = (3e^(x+c) - 1)^(1/3) y(0) = 1 1 = (3e^c -1)^(1/3) 1 = 3e^c - 1 2/3 = e^c c = ln(2/3) y = (3e^(x+ln(2/3)) - 1)^(1/3)

OpenStudy (anonymous):

hope i havent done any mistakes here though

OpenStudy (anonymous):

\[\frac{ dy }{ dx } = \frac{ (y^3+1) }{ y^2 } = \frac{ y^3 }{ y^2 } + \frac{ 1 }{ y^2 }\]\[\frac{ dy }{ dx } = y + y^-2\]\[y = \int\limits_{}^{}(y + y^-2) dx\]

OpenStudy (anonymous):

\[\frac{ DY }{ DX } <-- cannot divide by zero\]

OpenStudy (anonymous):

@Coolsector answer is like [2e^(3x)-1]^(1/3)

OpenStudy (anonymous):

let me check again

OpenStudy (anonymous):

(3e^(x+ln(2/3)) - 1)^(1/3) = (3e^(x) * e^(ln(2/3)) - 1)^(1/3) = (3e^(x) *(2/3)) - 1)^(1/3) = (2e^(x) - 1)^(1/3)

OpenStudy (anonymous):

my answer is the same .. we just had to simplify it

OpenStudy (anonymous):

because e^(ln(2/3)) = 2/3 :)

OpenStudy (anonymous):

@Coolsector oh ok haha thanks :D

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

@Coolsector how come this.. du / (3u) = dx becomes this? x +C = (1/3) ln|u| ? plz explain thanks :)

OpenStudy (anonymous):

@gaara you dont know any calculus do you?

OpenStudy (dumbcow):

@smoothieoeek \[\int\limits_{}^{}\frac{du}{3u} = \frac{1}{3}\int\limits_{}^{}\frac{du}{u} = \frac{1}{3} \ln |u|\] and \[\int\limits_{}^{}dx = x+C\]